# Learning with 3D rotations, a hitchhiker’s guide to SO(3)

### v1

@@ C.5 How does distance picking and computing distances in $\mathrm{SO}(3)$ affect rotation estimation?
C.6 Additional aspects on learning with rotations
@@
C.7 How does the ratio between distances induced by $g(r)$ look like for axis-angles and exponential coordinates?
@@
C.8 How does $\mathbb{R}^{9}$+SVD and $\mathbb{R}^{6}$+GSO compare in terms of computation times?
@@
Appendix D Further experiment details
@@
D.1 Details of Experiment 1 (Rotation estimation): Rotation from point clouds
@@
D.2 Details of Experiment 2: Cube image to / from rotation
@@
D.3 Details of Experiment 3 (Rotation estimation): 6D object pose estimation from RGB-D images
@@
D.4 Details of Experiment 4 (Feature prediction): $\mathrm{SO}(3)$ as input to Fourier series
The network is given by a 3-layer MLP with input size being determined by the rotation representation, hidden dimension $[256,256]$, ReLU activation functions, and the output size of one. The networks were trained for 400 Epochs using Adam (standard settings in PyTorch with a starting learning rate of $0.001$). At each iteration a batch of 64 data points was drawn from 800 train data points and validated on 400 validation points. After training, the best model with respect to the validation loss was used to compute a test loss on 1000 data points. For train, validation, and test loss we used the RMSE.
D.5 Experiment 5 (Rotation estimation): Inverse Kinematics with CMU MoCap data

### v2

@@ D.4 How does distance picking and computing distances in $\mathrm{SO}(3)$ affect rotation estimation?
D.5 Additional aspects on learning with rotations
@@
D.6 How does the ratio between distances induced by $g(r)$ look like for axis-angles and exponential coordinates?
@@
D.7 How does $\mathbb{R}^{9}$+SVD and $\mathbb{R}^{6}$+GSO compare in terms of computation times?
@@
Appendix E Further experiment details
@@
E.1 Details of Experiment 1 (Rotation estimation): Rotation from point clouds
@@
E.2 Details of Experiment 2: Cube image to / from rotation
@@
E.3 Details of Experiment 3 (Rotation estimation): 6D object pose estimation from RGB-D images
@@
E.4 Details of Experiment 4 (Feature prediction): $\mathrm{SO}(3)$ as input to Fourier series
The network is given by a 3-layer MLP with input size being determined by the rotation representation, hidden dimension $[256,256]$, ReLU activation functions, and the output size of one. The networks were trained for 400 Epochs using Adam (standard settings in PyTorch with a starting learning rate of $0.001$). At each iteration a batch of 64 data points was drawn from 800 train data points and validated on 200 validation points. After training, the best model with respect to the validation loss was used to compute a test loss on 1000 data points. For train, validation, and test loss we used the RMSE.
E.5 Experiment 5 (Rotation estimation): Inverse Kinematics with CMU MoCap data